# Benchmarking LLMs’ Judgments with No Gold Standard

### v1

@@ Appendix B Omitted Proofs
(This step follows the condition of bounded KL-divergence in Proposition 3.1.) $\geq$ $\sum_{x_{H}}\Pr[X_{H}=x_{H}]\left(\sum_{y}\Pr[Y=y\mid X_{H}=x_{H}]\log\Pr[Y=y\mid X_{H}=x_{H}]-\epsilon\right)$
@@
(This step follows the fact that logarithm is a proper scoring rule [Hendrickson and Buehler, 1971].) $\leq$ $\sum_{x_{L}}\Pr[X_{L}=x_{L}]\sum_{y}\Pr[Y=y\mid X_{L}=x_{L}]\log\Pr[Y=y\mid X_{L}=x_{L}]$
@@ Appendix C Implementation Details
In this section, we provide a detailed interpretation of how we implement our proposed metrics to facilitate the replication of our results. This includes comprehensive descriptions of all prompts utilized for generating peer-review judgments, preprocessing the judgments, and predicting the judgments.
C.1. Judgments Preprocessing
@@ C.2. Judgments Prediction
C.3. Peer-review Judgments Generation
In GRE-bench, we employ academic peer review as the task for candidate LLMs to generate informative judgments. Below is the prompt utilized to request each candidate LLM to produce its review judgments.
@@ C.3. Peer-review Judgments Generation
C.4. Summary of Author-stated Strengths and Weaknesses
@@ Appendix D Experimental Details
D.1. Degradations
D.1.1. Review Degradation: Sentence Deletion
@@
D.1.2. Review Degradation: Deletion & Completion
@@
D.1.3. Review Degradation: Abstract-only Review
@@
D.2. Manipulations
D.2.1. Review Manipulation: LLM Rephrasing
@@
D.2.2. Review Manipulation: Meaningless Elongation
@@
D.3. Baseline Metrics
In the paper, we use BLEU, ROUGE-L, BERTScore and LMExaminer as the baseline metrics. BLEU and ROUGE-L are both detail free. We employ NLTK as the tokenizer and utilized existing library functions to compute the BLEU and ROUGE-L scores. When calculating BERTScore, it is also necessary to split the text into several segments and embed each segment before computing the cosine similarity. We choose to split the text by sentence and then embedding each sentence individually. The model used for sentence embedding is stella_en_v5.
LMExaminer utilizes a specific LLM to score answers to questions. The original LMExaminer requires the input of the question, reference answer, and the answer to be scored into the LLM. However, since we are dealing with the subjective nature of peer review, we input the full text of the paper along with the review to be judged into the LLM and prompt it to score based on multiple dimensions. We employ GPT-4o as the examiner, and below are our prompts.
System Prompt
To judge the quality of this review, consider the following criteria:
Present your evaluation in the following format:
<analysis>
[Your detailed analysis here]
</analysis>
overall_assessment>
[Your overall assessment here]
</overall_assessment>
overall_score>
[Your final quality score here]
</overall_score>
User Prompt
[Paper]
{Full paper in Text Format}
[Review]
{Original Review Judgments}
@@ E.1. Implementation
model_config”: { base_model”:”unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit”, # The base model max_seq_length”: 4096, # The maximum sequence length load_in_4bit”: True, # Load the model in 4-bit }, lora_config”: { r”: 16, # The number of LoRA layers 8, 16, 32, 64 lora_alpha”:16, # The alpha value for LoRA lora_dropout”:0, # The dropout value for LoRA }, training_config”: { per_device_train_batch_size”: 2, # The batch size gradient_accumulation_steps”: 4, # The gradient accumulation steps warmup_steps”: 5, # The warmup steps max_steps”:0, # The maximum steps num_train_epochs”: 3, # The number of training epochs learning_rate”: 2e-4, # The learning rate optim :”adamw_8bit”, # The optimizer weight_decay : 0.01, # The weight decay lr_scheduler_type”: linear”, # The learning rate scheduler seed : 42, # The seed }
E.2. Experimental Results
@@
Table 8. Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation.
@@
Table 9. Standardized Mean Differences (SMD) of scores $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ after manipulations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score increase ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metric are not robust against the manipulation.
E.3. GRE-bench based on finetuned GEM and GEM-S
Figure 6. Results of GRE-bench based on finetuned GEM and GEM-S with 90% confidence intervals vs. model parameter size. The grey line represents the average human baseline, with the 90% confidence interval shaded in grey.

### v2

@@ Appendix B Omitted Proofs
$\geq$ $\sum_{x_{H}}\Pr[X_{H}=x_{H}]\left(\sum_{y}\Pr[Y=y\mid X_{H}=x_{H}]\log\Pr[Y=y\mid X_{H}=x_{H}]-\epsilon\right)$ (This step follows the condition of bounded KL-divergence in Proposition 3.1.)
@@
$\leq$ $\sum_{x_{L}}\Pr[X_{L}=x_{L}]\sum_{y}\Pr[Y=y\mid X_{L}=x_{L}]\log\Pr[Y=y\mid X_{L}=x_{L}]$ (This step follows the fact that logarithm is a proper scoring rule (Hendrickson & Buehler, 1971).)
@@ Appendix C Implementation Details
In this section, we provide a detailed interpretation of how we implement our proposed metrics to facilitate the replication of our results. This includes comprehensive descriptions of all prompts utilized for generating peer-review judgments, preprocessing the judgments, predicting the judgments, as well as other details that help reproduce the implementation. In all LLM API calls in our experiments, for reproducibility, we keep the temperature at 0 and the maximum (output) token count at 4000, unless otherwise specified.
C.1 Judgments Preprocessing
@@ C.2 Judgments Prediction
C.3 Peer-review Judgments Generation in GRE-bench
We provide the implementation details of GRE-bench where we employ academic peer review as the task to evaluate LLMs’ ability to generate informative judgments. Following Liang et al. (2023)’s research about LLM-generated peer review, we use the Sciencebeam pdf parser to convert the ICLR submissions in pdf format to text format, and removed the references and all appendices. When using the sciencebeam PDF parser, it retains all image and table titles and captions while disregarding other content. While a few ICLR submissions require OCR before parsing into text (5 out of all submissions in 2023), none of the 300 papers we experimented with required OCR, ensuring no OCR errors. Additionally, we have verified that the parsed text of papers in our dataset does not exceed the input token limit of the LLM.
While we acknowledge that PDF parsing is an inherently imperfect process that may occasionally introduce artifacts or miss certain elements, our experimental design mitigates potential biases because all LLM models in our evaluation receive identical parsed versions of each paper, and any parsing inconsistencies would affect all models equally, making the relative performance comparisons robust to such noise.
Below is the prompt utilized to request each candidate LLM to produce its review judgments.
@@ C.3 Peer-review Judgments Generation in GRE-bench
C.4 Summary of Author-stated Strengths and Weaknesses
@@ Appendix D Experimental Details
D.1 Baseline Metrics
In the paper, we use BLEU, ROUGE-L, BERTScore and LMExaminer as the baseline metrics. BLEU and ROUGE-L are both detail free. We employ NLTK as the tokenizer and utilized existing library functions to compute the BLEU and ROUGE-L scores.
We input the full text of the paper along with the review to be judged into the LLM and prompt it to score based on RQIs. We show our prompts below.
System Prompt
To judge the quality of this review, consider the following criteria:
Present your evaluation in the following format:
<analysis>
[Your detailed analysis here]
</analysis>
overall_assessment>
[Your overall assessment here]
</overall_assessment>
overall_score>
[Your final quality score here]
</overall_score>
User Prompt
[Paper]
{Full paper in Text Format}
[Review]
{Original Review Judgments}
D.2 Validation Workflow for Degradations and Manipulations
Output: Statistics metrics.
for $i=1$ to $n$ do
       Compute $s_{i}:=f(w_{i},x_{i},y_{i})$;
       Compute $s_{i}^{\prime}:=f(w_{i},x_{i}^{\prime},y_{i})$;
end for
ALGORITHM 2 Validation Workflow
D.3 Degradations
D.3.1 Review Degradation: Sentence Deletion
@@
D.3.2 Review Degradation: Deletion & Completion
@@
D.3.3 Review Degradation: Abstract-only Review
@@
D.4 Manipulations
D.4.1 Review Manipulation: LLM Rephrasing
@@
D.4.2 Review Manipulation: Meaningless Elongation
@@ E.1 Implementation
"model_config": { "base_model":"unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit", # The base model "max_seq_length": 4096, # The maximum sequence length "load_in_4bit": True, # Load the model in 4-bit }, "lora_config": { "r": 16, # The number of LoRA layers 8, 16, 32, 64 "lora_alpha":16, # The alpha value for LoRA "lora_dropout":0, # The dropout value for LoRA }, "training_config": { "per_device_train_batch_size": 2, # The batch size "gradient_accumulation_steps": 4, # The gradient accumulation steps "warmup_steps": 5, # The warmup steps "max_steps":0, # The maximum steps "num_train_epochs": 3, # The number of training epochs "learning_rate": 2e-4, # The learning rate "optim" :"adamw_8bit", # The optimizer "weight_decay" : 0.01, # The weight decay "lr_scheduler_type": "linear", # The learning rate scheduler "seed" : 42, # The seed }
E.2 Experimental Results
@@
Table 14: Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation.
@@
Table 15: Standardized Mean Differences (SMD) of scores $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ after manipulations of various evaluation metrics with 95% confidence intervals in parentheses. Significant score increase ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metric are not robust against the manipulation.
E.3 GRE-bench based on finetuned GEM and GEM-S
Figure 6: Results of GRE-bench based on finetuned GEM and GEM-S with 90% confidence intervals vs. model parameter size. The grey line represents the average human baseline, with the 90% confidence interval shaded in grey.